# A Nonparametric Bayesian Solution of the Empirical Stochastic Inverse Problem

### v1

@@ 4 Examples
Figure 11: Heatmaps of marginal distributions for the solution of the eSIP computed using a parametric fit to observed data obtained by dropping volley ball and tennis ball data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots. In the first and second plots, we indicate the intersection of $\Lambda$ with the highly probable range of $H_{0}$ given the realistic range of $g$ as a transparent horizontal rectangle. In the first plot, we indicate the intersection of $\Lambda$ with the highly probable range of $V_{0}$ as a transparent vertical rectangle.
@@
Next, we generate “noisy” output data by adding $70,000$ iid samples from the $\mathrm{Beta}(8,8)$ distribution scaled and shifted to $q_{i}-.35,q_{i}+.35]$ for each $q_{i}$ value, excepting the volleyball and tennis ball data, in Table 2. The other computational parameters remain the same. We show the solution in Fig. 12.
@@
Figure 12: Heatmaps of marginal distributions for the solution of the eSIP computed using noisy data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots. In the first and second plots, we indicate the intersection of $\Lambda$ with the highly probable range of $H_{0}$ given the realistic range of $g$ as a transparent horizontal rectangle. In the first plot, we indicate the intersection of $\Lambda$ with the highly probable range of $V_{0}$ as a transparent vertical rectangle.
@@
Interestingly, the eSIP solutions for the two different approaches to generating data are qualitatively similar.
@@
The solutions obtained by eliminating the data for the volleyball and the tennis ball are quite different than the solutions obtained for the full set of data. Notably, the structure in the densities imparted from the linear contours combined with the uniform prior is evident. A much larger set of possible parameter values produce solutions consistent with the observed data with relatively high probability.
@@
In particular, there is a large range of $H_{0}$ and $V_{0}$ values that are consistent with $g$ close to $9.8$. We plot the event for $g\in[9.78,9.82]$ in Fig. 11 and 12 as vertical rectangles. Using the intersection of that event with the region of relatively high probability determines ranges for $H_{0}$ of $[37.5,42.5]$ respectively $[36.75,43]$ of relatively high probability. Lastly, this indicates that $V_{0}$ is in the range of $[0,1]$. We also see that the region of relatively high probability has the property that larger values of $H_{0}$ are associated with more negative values of $V_{0}$ and vice versa, which fits intuition.
@@
Solving the eSIP for the bowling ball data with an informed prior.
@@
We show the solution in Fig. 13. The effect of using a prior with a small effective support is obvious as the solution is also concentrated in a small region. The mode of the solution is near $(35.7\;.5\;9.9)^{\top}$ and the event of relatively high probability places $g$ within $[9.78,9.95]$.
@@
Figure 13: Heatmaps of marginal distributions for the solution of the eSIP computed using noisy data for the bowling ball and an informed prior. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$.
@@ Conclusions.
Including the volleyball and tennis ball data in the eSIP for the simplified model results in solutions that compensate for neglecting drag by putting higher probability on combinations of the parameters that produce longer drop times. Eliminating those data results in a solution that more accurately reflects the structure imparted by the model. Using an informed prior for the eSIP with the bowling ball data lead to a solution that is locally concentrated near a point, but did not yield a more accurate estimate.
5 An accept-reject solution method for the eSIP
As a partial connection to Markov Chain Monte Carlo methods typically used for Bayesian statistics, we adapt an accept-reject algorithm from [16] to construct a method for the eSIP that produces a collection of independent random samples approximately distributed according to the posterior distribution. The accept-reject criteria is based on the “update” weight,
@@ 8     Generate a random sample $\xi$ from the uniform distribution on $[0,1]$
Algorithm 1 Accept-Reject Algorithm for the eSIP
The samples $\{\breve{\lambda}_{k}\}_{k=1}^{\breve{N}}$ are approximately distributed according to the SIP solution distribution $P_{\Lambda}$ with density,
@@
In our experience, the solution method based on reweighting random samples is generally more computationally efficient, especially if there is a high reject rate in the accept-reject method.
@@
We apply the accept-reject algorithm 1 to solve the eSIP in Example 1 using synthetic data. We hold the discretization and computational parameters fixed as in Example 1 except as noted. We choose $K=100$ for a uniform partition of $\mathcal{D}=[0,1]$. We use $N=10^{6}$ samples from the data generating distribution $P_{\Lambda}^{\mathrm{d}}$ to compute the synthetic data used to construct $\hat{\rho}_{\mathcal{D}}$ and $N=10^{6}$ samples from uniform prior on $\Lambda$ to construct $\widehat{\tilde{\rho}}_{\mathrm{p},\mathcal{D}}$.
@@
Figure 14: Left: Plots of $\hat{\rho}_{\mathcal{D}}$ (blue) and $\widehat{\tilde{\rho}}_{\mathrm{p},\mathcal{D}}$ (red). Center: Scatter plots of samples from the data generating distribution (red), uniform prior (green), and the posterior distribution (blue). We see that the accepted samples lie in the “shadow” of the data generating distribution along the contours. Right: Scatter plot of uniform prior samples accepted by the accept-reject algorithm that are distributed according to the posterior distribution on top of the heatmap of the empirical posterior computed from re-weighting samples distributed according to the prior.
We emphasize that the importance sampling and accept-reject methodologies are two different approaches for estimating the same solution.
@@ 6 Conclusion
In this paper, we consider the problem of solving an inverse problem in which a complex system is described by a computer model and system behavior is governed by parameters in the model but it is impossible to observe the values of the input parameters. So, they must be inferred using model evaluations and observations from the system. We formulate and solve the empirical Stochastic Inverse Problem, which is to determine a probability distribution for the parameters governing the system.
We develop and analyze a nonparametric Bayesian approach to defining a unique posterior distribution for the parameters given a prior. We prove that the solution exists and give a concrete expression for the solution using disintegration of measures. Conditions where the solution can be written in terms of conditional densities is developed, and we prove that under general conditions, the solution is continuous a.e. We also show that the solution corresponding to the uniform prior has maximum entropy. A nonparametric solution method based on random sampling is proposed, with an analysis of its accuracy and convergence properties, along with an alternative accept–reject method. Finally, we illustrate with a number of examples, including one that deals with the situation in which there is insufficient data for a nonparametric approach and a simulation-based inference approach is used.
Future work includes incorporating random effects to broaden the range of applications for the SIP framework, including linking the eSIP to classic Bayesian statistics and formulating and solving the eSIP for multiple experiments on one system. Another direction is the development of more comprehensive methods for situations where the available data is limited and simulation-based is required. Ongoing efforts also focus on extending SIP solutions to infinite-dimensional spaces and investigating the use of Markov Chain Monte Carlo methods and machine learning for the solution of the eSIP.
@@ Appendix A Proofs
Let ${P}_{\Lambda}$ be the solution of the SIP corresponding any prior $P_{\mathrm{p}}$ and let $\overline{P}_{\Lambda}$ denote the solution corresponding to the uniform prior. By Theorem 2.9, these have respective densities,
@@
For simplicity of notation, we prove the result for the uniform prior. The general result follows using the same proof with minor alterations. We first prove a special case.
@@
Without loss of generality, we assume $i_{1}=1,i_{2}=2,\dots,i_{m}=m$. We let $\lambda^{(m)}=(\lambda_{1},\dots,\lambda_{m})$, $\hat{\lambda}^{(m)}=(\lambda_{m+1},\dots,\lambda_{n})$, so $\lambda$ can be formally written as $\lambda\sim(\lambda^{(m)},\hat{\lambda}^{(m)})$. We define $\pi_{\hat{m}}:\mathbb{R}^{n}\to\mathbb{R}^{n-m}$ as $\pi_{\hat{m}}(\lambda)=\hat{\lambda}^{(m)}$.
For a $\lambda_{0}\in\Lambda\cap Q^{-1}(q_{0})$, since $J_{Q}^{(m)}(\lambda_{0})$ is full rank, Theorem A.1 implies that $Q(\lambda)-q=0$ determines a unique continuous function $h:N(\hat{\lambda}_{0}^{(m)},r)\times N(q_{0},r)\to N(\lambda_{0}^{(m)},r)$. Without loss of generality we assume $h$ is defined on $\overline{N(\hat{\lambda}_{0}^{(m)},r)}\times\overline{N(q_{0},r)}$. Since the collection $\{N(\hat{\lambda}^{(m)},r)\times N(\lambda^{(m)},r)\}$ covers the compact set $\Lambda\cap Q^{-1}(q_{0})$, there is a finite subcover $\{N(\hat{\lambda}_{k}^{(m)},r_{k})\times N(\lambda_{k}^{(m)},r_{k})\},k=1,\dots,K$. Let $r_{0}=\min_{k}r_{k}$. There is a unique continuous function $h:\left(\bigcup_{k=1}^{K}\overline{N(\hat{\lambda}_{k}^{(m)},r_{k})}\right)\times\overline{N(q_{0},r_{0})}\to\bigcup_{k=1}^{K}N(\lambda_{k}^{(m)},r_{k})$. We have
$\tilde{\rho}_{\mathcal{D}}(q_{0})=\int_{\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))}\frac{1}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q_{0}),\hat{\lambda}^{(m)})\right|}\,d\hat{\lambda}^{(m)}.$
Since $J_{Q}^{(m)}(\lambda)$ is a continuous function, $\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},y),\hat{\lambda}^{(m)})\right|^{-1}$ is continuous on a compact domain $\left(\bigcup_{k=1}^{K}\overline{N(\hat{\lambda}_{k}^{(m)},r_{k})}\right)\times\overline{N(q_{0},r_{0})}$, which means it is uniformly continuous and bounded. Let $M>0$ be bound so $\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},y),\hat{\lambda}^{(m)})\right|^{-1}<M$.
@@
Define $\lambda^{(m+1)}=(\lambda^{(m)},\lambda_{j}),\hat{\lambda}^{(m+1)}=\hat{\lambda}^{(m)}\setminus\lambda_{j}$, where we write $\lambda\sim(\lambda^{(m+1)},\hat{\lambda}^{(m+1)})$. By Theorem A.1, there exists a unique continuous function,
@@
which satisfies $Q^{*}(g(\hat{\lambda}^{(m+1)},y),\lambda^{(m)},\lambda_{j})=0$. Without loss of generality, we assume $g$ is defined on $\overline{N(\hat{\lambda}_{0}^{(m+1)},r^{\prime})}\times\overline{N(q_{0},r^{\prime})}$. By compactness of the domain, $g$ is uniformly continuous. The set $\{N(\hat{\lambda}^{(m+1)},r^{\prime})\times N(\lambda^{(m)},r^{\prime})\times N(\lambda_{j},r^{\prime})\}$ covers $\partial\Lambda\cap Q^{-1}(y)$ for $y\in N(q_{0},r^{\prime})$. We assume $r^{\prime}$ is sufficiently small so that
@@
Since $\partial\Lambda\cap Q^{-1}(q_{0})$ is compact, there is a finite subcover $\{N(\hat{\lambda}_{\ell}^{(m+1)},r_{\ell}^{\prime})\times N(\lambda_{\ell}^{(m)},r_{\ell}^{\prime})\times N(\lambda_{\ell j},r_{\ell}^{\prime})\}$, $\ell=1,\dots,L$. Let $r_{0}^{\prime}=\min\{r^{\prime}_{\ell},\ell=1,\cdots,L;r_{0}\}$ then for any $y\in N(q_{0},r_{0}^{\prime})$, the sets also cover $\partial\Lambda\cap Q^{-1}(y)$. So Condition 2 holds for all $y\in N(q_{0},r_{0}^{\prime})$. Thus
@@
is covered by $\{N(\hat{\lambda}_{k}^{(m)},r_{k})\times N(\lambda_{k}^{(m)},r_{k})\times N(q_{0},r_{0})\}$.
As a result, $\{N(\hat{\lambda}_{k}^{(m)},r_{k})\times N(\lambda_{k}^{(m)},r_{k})\}$ covers $\Lambda\cap Q^{-1}(y)$ for all $y\in N(q_{0},r_{0}^{\prime})$. Thus,
$\tilde{\rho}_{\mathcal{D}}(y)=\int_{\pi_{\hat{m}}(\Lambda\cap Q^{-1}(y))}\frac{1}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},y),\hat{\lambda}^{(m)})\right|}\,d\hat{\lambda}^{(m)}.$
For $y\in N(q_{0},r^{\prime}_{0})$, let $C=\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))\triangle\pi_{\hat{m}}(\Lambda\cap Q^{-1}(y))$. Thus,
@@
$\qquad+\int_{C}M\,d\hat{\lambda}^{(m)}.$
For the first term, uniform continuity implies that given $\epsilon>0$, there exists $r_{\epsilon}<r_{0}^{\prime}$ such that for $y\in N(q_{0},r_{\epsilon})$, and $\hat{\lambda}^{(m)}\in\left(\bigcup_{k=1}^{K}N(\hat{\lambda}_{k}^{(m)},r_{k})\right)$,
@@
For the second term, we need $\hat{\mu}^{(m)}(B)$ to be small when $q$ is close to $q_{0}$. Since $\{\pi_{\hat{m}}\big(N(\hat{\lambda}_{\ell}^{(m+1)},$ $r_{\ell}^{\prime})\times N(\lambda_{\ell}^{(m)},r_{\ell}^{\prime})\times N(\lambda_{\ell j},r_{\ell}^{\prime})\big)=N(\hat{\lambda}_{\ell}^{(m+1)},r_{\ell}^{\prime})\times N(\lambda_{\ell j},r_{\ell}^{\prime})\}$ covers $\pi_{\hat{m}}(\partial\Lambda\cap Q^{-1}(q))$ when $q\in N(q_{0},r_{0}^{\prime})$, then
$\hat{\mu}^{(m)}(B)=\hat{\mu}^{(m)}(\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))\triangle\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q)))$
@@
$\leq\int_{N(\hat{\lambda}_{\ell}^{(m+1)},r_{\ell}^{\prime})}\left|g(\hat{\lambda}_{\ell}^{(m+1)},q_{0})-g(\hat{\lambda}_{\ell}^{(m+1)},q)\right|\,d\hat{\lambda}^{(m+1)}.$
By the uniform continuity of $g$, there exists an $r_{\epsilon}^{\prime}\leq r_{0}^{\prime}$ such that for all $q\in N(q_{0},r_{\epsilon}^{\prime})$, $\hat{\mu}^{(m)}(B)<\epsilon$. Finally, we have for all $y\in N(q_{0},r)$, $r=\min(r_{\epsilon},r_{\epsilon}^{\prime})$,
$|\tilde{\rho}_{\mathcal{D}}(q_{0})$ $-\tilde{\rho}_{\mathcal{D}}(y)|$
@@
So $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous at $q_{0}$. Furthermore, the two conditions hold for all $q$ in $N(q_{0},r_{0}^{\prime})$, so $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous in $N(q_{0},r)$ as well. ∎

### v3

@@ 4 Examples
Figure 12: Heatmaps of marginal distributions for the posterior computed using a parametric fit to observed data obtained by dropping volley ball and tennis ball data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots. In the first and second plots, we indicate the intersection of $\Lambda$ with the highly probable range of $H_{0}$ given the realistic range of $g$ as a transparent horizontal rectangle. In the first plot, we indicate the intersection of $\Lambda$ with the highly probable range of $V_{0}$ as a transparent vertical rectangle.
@@
Next, we generate “noisy” output data by adding $70,000$ iid samples from the $\mathrm{Beta}(8,8)$ distribution scaled and shifted to $q_{i}-.35,q_{i}+.35]$ for each $q_{i}$ value, excepting the volleyball and tennis ball data, in Table 2. The other computational parameters remain the same. We show the posterior in Fig. 13.
@@
Figure 13: Heatmaps of marginal distributions for the posterior computed using noisy data. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$. We indicate the part of $\Lambda$ intersecting the realistic range of uncertainty for $g$ near sea level as a transparent vertical rectangle in the second and third plots. In the first and second plots, we indicate the intersection of $\Lambda$ with the highly probable range of $H_{0}$ given the realistic range of $g$ as a transparent horizontal rectangle. In the first plot, we indicate the intersection of $\Lambda$ with the highly probable range of $V_{0}$ as a transparent vertical rectangle.
@@
Interestingly, the posteriors for the two different approaches to generating data are qualitatively similar.
@@
The posterior obtained by eliminating the data for the volleyball and the tennis ball are quite different than the posteriors obtained for the full set of data. Notably, the structure in the densities imparted from the linear contours combined with the uniform prior is evident. A much larger set of possible parameter values produce computer model solutions consistent with the observed data with relatively high probability.
@@
In particular, there is a large range of $H_{0}$ and $V_{0}$ values that are consistent with $g$ close to $9.8$. We plot the event for $g\in[9.78,9.82]$ in Fig. 12 and 13 as vertical rectangles. Using the intersection of that event with the region of relatively high probability determines ranges for $H_{0}$ of $[37.5,42.5]$ respectively $[36.75,43]$ of relatively high probability. Lastly, this indicates that $V_{0}$ is in the range of $[0,1]$. We also see that the region of relatively high probability has the property that larger values of $H_{0}$ are associated with more negative values of $V_{0}$ and vice versa, which fits intuition.
@@
Calibrating the computer model using bowling ball data and an informed prior.
@@
We show the posterior in Fig. 14. The effect of using a prior with a small effective support is obvious as the posterior is also concentrated in a small region. The mode of the posterior is near $(35.7\;.5\;9.9)^{\top}$ and the event of relatively high probability places $g$ within $[9.78,9.95]$.
@@
Figure 14: Heatmaps of marginal distributions for the posterior computed using noisy data for the bowling ball and an informed prior. Left: $H_{0}$ vs $V_{0}$. Center: $H_{0}$ vs $g$. Right: $V_{0}$ vs $g$.
@@ Conclusions.
Including the volleyball and tennis ball data results in posteriors that appear to “compensate” for neglecting drag by putting higher probability on combinations of the parameters that produce longer drop times. Eliminating those data results in posteriors that more accurately reflects the structure imparted by the model. Using an informed prior for the eSIP with the bowling ball data lead to a posterior that is locally concentrated near a point, but did not yield a more accurate estimate.
5 An estimator based on accept-reject sampling
As a partial connection to Markov Chain Monte Carlo methods typically used for Bayesian statistics, we adapt a methodology from [16] to construct an estimator that produces a collection of independent random samples approximately distributed according to the posterior. This is performed using accept-reject using a criteria based on the “update” weight,
@@ 8     Generate a random sample $\xi$ from the uniform distribution on $[0,1]$
Algorithm 1 Accept-Reject Algorithm
The samples $\{\breve{\lambda}_{k}\}_{k=1}^{\breve{N}}$ are approximately distributed according to posterior $P_{\Lambda}$ with density,
@@
In our experience, the estimator based on reweighting random samples is generally more computationally efficient, especially if there is a high reject rate in the accept-reject estimator.
@@
We apply the accept-reject algorithm 1 to Example 1 using synthetic data. We hold the discretization and computational parameters fixed as in Example 1 except as noted. We choose $K=100$ for a uniform partition of $\mathcal{D}=[0,1]$. We use $N=10^{6}$ samples from the data generating distribution $P_{\Lambda}^{\mathrm{d}}$ to compute the synthetic data used to construct $\hat{\rho}_{\mathcal{D}}$ and $N=10^{6}$ samples from uniform prior on $\Lambda$ to construct $\widehat{\tilde{\rho}}_{\mathrm{p},\mathcal{D}}$.
@@
Figure 15: Left: Plots of $\hat{\rho}_{\mathcal{D}}$ (blue) and $\widehat{\tilde{\rho}}_{\mathrm{p},\mathcal{D}}$ (red). Center: Scatter plots of samples from the data generating distribution (red), uniform prior (green), and the posterior (blue). We see that the accepted samples lie in the “shadow” of the data generating distribution along the contours. Right: Scatter plot of uniform prior samples accepted by the accept-reject algorithm that are distributed according to the posterior on top of the heatmap of the empirical posterior computed from re-weighting samples distributed according to the prior.
We emphasize that the importance sampling and accept-reject methodologies are two different approaches for estimating the same posterior.
@@ 6 Conclusion
In this paper, we consider the problem of calibrating a computer model for a distribution on physical parameters. We formulate calibration as an empirical Stochastic Inverse Problem. We develop and analyze a nonparametric Bayesian methodology for estimating a unique posterior for a given prior. We prove that the posterior exists using disintegration of measures. We provide conditions under which the posterior can be represented by a conditional density. We prove that under general conditions, the posterior is continuous a.e. We also show that the posterior corresponding to the uniform prior has maximum entropy. A nonparametric estimator based on random sampling is constructed and analyzed along with an alternative accept–reject estimator. Finally, we illustrate with a number of examples, including one that deals with the situation in which there is insufficient data for a nonparametric approach and a simulation-based inference approach is used.
Future work includes incorporating random effects to broaden the range of applications for the SIP framework, including more closely linking the eSIP to Bayesian statistics and calibration of computer models using data from multiple experiments. Another direction is the development of more comprehensive methods for situations where the available data is limited and simulation-based is required. Ongoing efforts also focus on extending this calibration approach to infinite-dimensional spaces and investigating the use of Markov Chain Monte Carlo methods and machine learning for estimating the posterior.
@@ Appendix A Proofs
Let ${P}_{\Lambda}$ be the posterior corresponding any prior $P_{\mathrm{p}}$ and let $\overline{P}_{\Lambda}$ denote the posterior corresponding to the uniform prior. By Theorem 2.9, these have respective densities,
@@
For simplicity of notation, we first prove the result for the uniform prior. We treat the general case afterwards.
@@
We first show the result for the uniform prior. Without loss of generality, we assume $i_{1}=1,i_{2}=2,\dots,i_{m}=m$. We let $\lambda^{(m)}=(\lambda_{1},\dots,\lambda_{m})$, $\hat{\lambda}^{(m)}=(\lambda_{m+1},\dots,\lambda_{n})$, so $\lambda$ can be formally written as $\lambda\sim(\lambda^{(m)},\hat{\lambda}^{(m)})$. We define $\pi_{\hat{m}}:\mathbb{R}^{n}\to\mathbb{R}^{n-m}$ as $\pi_{\hat{m}}(\lambda)=\hat{\lambda}^{(m)}$.
For a $\lambda_{0}\in\Lambda\cap Q^{-1}(q_{0})$, since $J_{Q}^{(m)}(\lambda_{0})$ is full rank, Theorem A.1 implies that $Q(\lambda)-q=0$ determines a unique continuous function $h:N(\hat{\lambda}_{0}^{(m)},r)\times N(q_{0},r)\to N(\lambda_{0}^{(m)},r)$. Without loss of generality we assume $h$ is defined on $\overline{N(\hat{\lambda}_{0}^{(m)},r)}\times\overline{N(q_{0},r)}$. Since the collection $\{N(\hat{\lambda}^{(m)},r)\times N(\lambda^{(m)},r):\lambda\in\Lambda\cap Q^{-1}(q_{0})\}$ covers the compact set $\Lambda\cap Q^{-1}(q_{0})$, there is a finite subcover $\{N(\hat{\lambda}_{k}^{(m)},r_{k})\times N(\lambda_{k}^{(m)},r_{k})\},k=1,\dots,K$. Let $r_{0}=\min_{k}r_{k}$. There is a unique continuous function $h:\left(\bigcup_{k=1}^{K}\overline{N(\hat{\lambda}_{k}^{(m)},r_{k})}\right)\times\overline{N(q_{0},r_{0})}\to\bigcup_{k=1}^{K}N(\lambda_{k}^{(m)},r_{k})$. We have
$\tilde{\rho}_{\mathcal{D}}(q_{0})=\int_{\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))}\frac{1}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q_{0}),\hat{\lambda}^{(m)})\right|}\,d\hat{\lambda}^{(m)},$
where we use $|J_{Q}|$ to denote $|\det J_{Q}|$ to simplify notation. Since $J_{Q}^{(m)}(\lambda)$ is a continuous function, $\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})\right|^{-1}$ is continuous on a compact domain $\left(\bigcup_{k=1}^{K}\overline{N(\hat{\lambda}_{k}^{(m)},r_{k})}\right)\times\overline{N(q_{0},r_{0})}$, which means it is uniformly continuous and bounded. Let $M>0$ be bound so $\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})\right|^{-1}<M$.
@@
We define $\lambda^{(m+1)}=(\lambda^{(m)},\lambda_{j}),\hat{\lambda}^{(m+1)}=\hat{\lambda}^{(m)}\setminus\lambda_{j}$, where we write $\lambda\sim(\lambda^{(m+1)},\hat{\lambda}^{(m+1)})$. By Theorem A.1, there exists a unique continuous function,
@@
which satisfies $Q^{*}(g(\hat{\lambda}^{(m+1)},y),\hat{\lambda}^{(m+1)})=0$. Without loss of generality, we assume $g$ is defined on $\overline{N(\hat{\lambda}_{0}^{(m+1)},r^{\prime})}\times\overline{N(q_{0},r^{\prime})}$. By compactness of the domain, $g$ is uniformly continuous. The collection of sets $\{N(\hat{\lambda}^{(m+1)},r^{\prime})\times N(\lambda^{(m)},r^{\prime})\times N(\lambda_{j},r^{\prime}):\lambda\in\partial\Lambda\cap Q^{-1}(q_{0})\}$ covers $\partial\Lambda\cap Q^{-1}(y)$ for $y\in N(q_{0},r^{\prime})$. We assume $r^{\prime}$ is sufficiently small so that
@@
Since $\partial\Lambda\cap Q^{-1}(q_{0})$ is compact, there is a finite subcover $\{N(\hat{\lambda}_{\ell}^{(m+1)},r_{\ell}^{\prime})\times N(\lambda_{\ell}^{(m)},r_{\ell}^{\prime})\times N(\lambda_{\ell j},r_{\ell}^{\prime})\}$, $\ell=1,\dots,L$. Let $r_{0}^{\prime}=\min\{r^{\prime}_{\ell},\ell=1,\cdots,L;r_{0}\}$ then for any $q\in N(q_{0},r_{0}^{\prime})$, the sets also cover $\partial\Lambda\cap Q^{-1}(q)$. So Condition 2 holds for all $q\in N(q_{0},r_{0}^{\prime})$. Thus,
@@
is covered by $\{N(\hat{\lambda}_{k}^{(m)},r_{k})\times N(\lambda_{k}^{(m)},r_{k})\times N(q_{0},r_{0}):k=1,\cdots,K\}$.
As a result, $\{N(\hat{\lambda}_{k}^{(m)},r_{k})\times N(\lambda_{k}^{(m)},r_{k})\}$ covers $\Lambda\cap Q^{-1}(q)$ for all $q\in N(q_{0},r_{0}^{\prime})$. Thus,
$\tilde{\rho}_{\mathcal{D}}(q)=\int_{\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q))}\frac{1}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})\right|}\,d\hat{\lambda}^{(m)}.$
For $q\in N(q_{0},r^{\prime}_{0})$, let $C=\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))\triangle\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q))$. Thus,
@@
$\qquad+\int_{C}M\,d\hat{\lambda}^{(m)}.$ (A.4)
For the first term, uniform continuity implies that given $\epsilon>0$, there exists $r_{\epsilon}<r_{0}^{\prime}$ such that for $q\in N(q_{0},r_{\epsilon})$, and $\hat{\lambda}^{(m)}\in\left(\bigcup_{k=1}^{K}N(\hat{\lambda}_{k}^{(m)},r_{k})\right)$,
@@
For the second term, we need $\hat{\mu}^{(m)}(C)$ to be small when $q$ is close to $q_{0}$. Since $\{\pi_{\hat{m}}\big(N(\hat{\lambda}_{\ell}^{(m+1)},$ $r_{\ell}^{\prime})\times N(\lambda_{\ell}^{(m)},r_{\ell}^{\prime})\times N(\lambda_{\ell j},r_{\ell}^{\prime})\big)=N(\hat{\lambda}_{\ell}^{(m+1)},r_{\ell}^{\prime})\times N(\lambda_{\ell j},r_{\ell}^{\prime})\}$ covers $\pi_{\hat{m}}(\partial\Lambda\cap Q^{-1}(q))$ when $q\in N(q_{0},r_{0}^{\prime})$, then
$\hat{\mu}^{(m)}(C)=\hat{\mu}^{(m)}(\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))\triangle\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q)))$
@@
$\leq\int_{N(\hat{\lambda}_{\ell}^{(m+1)},r_{\ell}^{\prime})}\left|g(\hat{\lambda}^{(m+1)},q_{0})-g(\hat{\lambda}^{(m+1)},q)\right|\,d\hat{\lambda}^{(m+1)}.$
By the uniform continuity of $g$, there exists an $r_{\epsilon}^{\prime}\leq r_{0}^{\prime}$ such that for all $q\in N(q_{0},r_{\epsilon}^{\prime})$, $\hat{\mu}^{(m)}(B)<\epsilon$. Finally, we have for all $q\in N(q_{0},r)$, $r=\min(r_{\epsilon},r_{\epsilon}^{\prime})$,
$|\tilde{\rho}_{\mathcal{D}}(q_{0})$ $-\tilde{\rho}_{\mathcal{D}}(q)|$
@@
So $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous at $q_{0}$. Furthermore, the two conditions hold for all $q$ in $N(q_{0},r_{0}^{\prime})$, so $\tilde{\rho}_{\mathcal{D}}(q)$ is continuous in $N(q_{0},r)$ as well.
Next, we treat a general prior. Arguing as for (A.4), we have
$|\tilde{\rho}_{p,\mathcal{D}}(q_{0})$ $-\tilde{\rho}_{p,\mathcal{D}}(q)|$
$\leq\int_{\pi_{\hat{m}}(\Lambda\cap Q^{-1}(q_{0}))}\left|\frac{\rho_{p}(h(\hat{\lambda}^{(m)},q_{0}),\hat{\lambda}^{(m)})}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q_{0}),\hat{\lambda}^{(m)})\right|}-\frac{\rho_{p}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})\right|}\right|\,d\hat{\lambda}^{(m)}$
$\qquad+\int_{C}M_{2}\,d\hat{\lambda}^{(m)},$
where $M_{2}$ is a constant such that,
$\left|\frac{\rho_{p}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})}{\left|J_{Q}^{(m)}(h(\hat{\lambda}^{(m)},q),\hat{\lambda}^{(m)})\right|}\right|<M_{2}\textrm{ on }\bigcup_{k=1}^{K}\left(N(\hat{\lambda}_{k}^{(m)},r_{k})\right)\times N(q_{0},r_{0}).$
Using the same argument as for the case of a uniform prior, there exists a $\delta>0$ such that $|q-q_{0}|<\delta$ implies $|\tilde{\rho}_{p,\mathcal{D}}(q_{0})-\tilde{\rho}_{p,\mathcal{D}}(q)|<\epsilon$.